A heat exchanger has been designed for use in a chemical process. A standard type of heat exchanger with negligible scrap value costs 4000s and will have a useful life of 6 years. Another proposed heat exchanger of equivalent design capacity costs 6800s but will have a useful life of 10 years and a scrap value of 800S. Assuming an effective compound interest of 8% percent per year, determine which heat exchanger is cheaper by comparing the capitalized costs?
Answer #1

Annual cost of first hear exchanger = 4000*(A/P,8%,6)

= 4000*0.216315

= 865.26

Capitalized cost of first heat exchanger = 865.26 / 0.08 = 10815.75

Annual cost of second hear exchanger = 6000*(A/P,8%,10) - 800*(A/F,8%,10)

= 6000*0.149029 - 800*0.069029

= 838.95

Capitalized cost of second heat exchanger = 838.95 / 0.08 = 10486.89

As Capitalized cost of second heat exchanger is less, it should be selected